Determine the values of the variable for which the expression is defined as a real number. (Enter your answer using interval notation.)
square root of x^2 − 16

I squared the entire thing, so that took care of the square root. Then I factored it into (x-4) (x+4) but thats as far as I got...
help, please? :(

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Determine the values of the variable for which the expression is defined as a real number. (Enter your answer using interval notation.)
square root of x^2 − 16
------
sqrt(x^2-16)
----
For the value to be a Real Number, x^2-16 must be >= 0
-----
Solve:
x^2-16 >= 0
(x-4)(x+4) >= 0
------
Ans: x >= 4 or x <= -4
-------------
Cheers,
Stan H.
------------

Answer by ikleyn(52867)

.
Many similar examples are considered in the lesson  Domain of a function which involves a quadratic polynomial under the square root operator  in this site.

It is free of charge and without registration.